OpenStudy (anonymous):

solve with substitution x+y=461 6x+5y=2530

OpenStudy (anonymous):

1) For the first equation x+y=461, subtract x to both sides to get y=461-x 2) You know that y=461-x now 3) Since that is your y value, plug it into the second equation 4) You should get 6x+5(461-x)=2530 5) Distribute 5 to 461 and -x to get an equation: 6x+2305+5x=2530 6) Combine like terms which are 6x and 5x to get 11x+2305=2530 7) Subtract 2305 from both sides to get 11x=225 8) Divide 11 from both sides to get 20 5/11 9) That is your x value and in order to find the y value plug in the x value to any equation 10) So, 20 5/11+y=461 11) Subtract 20 5/11 from both sides to get y= 440 6/11 12) The same y value should come for the second equation as well 13) you x value is 20 5/11 and y value is 440 6/11 Hope that helps :D